OpenStudy (anonymous):

[9.05] What are the approximate solutions of 2x2 + 5x = 4 rounded to the nearest hundredth?

OpenStudy (campbell_st):

rewrite the equation \[2x^2 + 5x - 4 = 0\] then using the general quadratic equation \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] where a = 2, b = 5 and c = -4 substitute and evaluate

OpenStudy (anonymous):

\[x=\frac{1}{4} \left(-5\pm \sqrt{57}\right) \]\[\{-3.14,0.64\} \]

OpenStudy (anonymous):

What are the exact solutions of x2 + 7x + 3 = 0?

OpenStudy (anonymous):

\[x=\frac{1}{2} \left(-7\pm \sqrt{37}\right) \]